OpenStudy (anonymous):

Find two consecutive integers whose sum is 113. Which of the following equations could be used to solve the problem? 2x = 113 2x + 1 = 113 2x + 2 = 113 x2 + 1= 113

OpenStudy (anonymous):

let x,x+1 be consecutive integers. x+x+1=113 ?

OpenStudy (anonymous):

what do you mean ? 2x+1=118 ?

OpenStudy (anonymous):

113*

OpenStudy (anonymous):

2x+1=113 is correct.